# Using Machine Learning to Explore Defect Configurations in Cd/Zn-Se/Te Compounds

### v1

@@ 9 Conclusions
CdSexTe1-x solar cells are fundamentally constrained by defect physics: deep-level nonradiative centers from native defects and impurities limit open-circuit voltage, dopants such as Cu and As often lead to unhelpful complexes, and extended defects at interfaces and grain boundaries act as sinks for charge and sites for defect clustering. While DFT remains the gold standard for resolving these mechanisms, its cost prevents exhaustive exploration across alloy compositions, charge states, and structural motifs. To overcome these barriers, we introduced the DeFecT-FF framework in this work, a crystal graph-based active learning-driven MLFF model trained on both semi-local and hybrid functional calculations for thousands of charged and neutral structures spanning the Cd/Zn–S/Se/Te bulk, alloy, interface, and grain boundary configurations with native and extrinsic defects and defect complexes. DeFecT-FF predicts energies and forces across charge states, enabling rapid geometry optimization and defect formation energy evaluation. We demonstrated the utility of these models by identifying low energy configurations of device-relevant defects and performing HSE06+SOC calculations to understand their energetics and defect levels.
In practice, the DeFecT-FF framework reduces single defect optimization time from $\sim$12–14 h (HSE06) to $\sim$1–2 min while retaining near-DFT accuracy, transforming comprehensive, composition- and charge-resolved defect surveys from intractable to routine. We have deployed this framework as part of a Jupyter notebook-based nanoHUB tool which will allow users to upload CIF files of Cd/Zn-Te/Se/S structures, auto-generate relevant defects or complexes, and compute their defect formation energies as functions of Fermi level and chemical potentials conditions, bypassing expensive first principles workflows. Together, these advances provide a scalable, charge-aware pathway to map defect landscapes in chemistries relevant to CdSeTe solar cell devices and beyond, accelerating the dopant/process optimization and ultimately closing the voltage deficit in this important thin-film photovoltaic platform.

### v2

@@ 5 Conclusions
CdSexTe1-x solar cells are fundamentally constrained by defect physics: deep-level nonradiative centers from native defects and impurities limit open-circuit voltage, dopants such as Cu and As often lead to unhelpful complexes, and extended defects at interfaces and grain boundaries act as sinks for charge and sites for defect clustering. While hybrid-functional DFT remains the gold standard for resolving these mechanisms, its cost prevents exhaustive exploration across alloy compositions, charge states, and structural motifs. To overcome these barriers, we introduced the DeFecT-FF framework in this work, a crystal graph-based active learning-driven MLFF model trained on data from both semi-local GGA and hybrid HSE06 calculations for thousands of charged and neutral structures spanning the Cd/Zn–S/Se/Te chemical space, with a wide variety of native and extrinsic defects and defect complexes considered. DeFecT-FF predicts energies and forces across charge states, enabling rapid geometry optimization and defect formation energy evaluation. We demonstrated the utility of these models by identifying low energy configurations of device-relevant defects and performing HSE06+SOC calculations to understand their energetics and defect levels.
In practice, the DeFecT-FF framework reduces single defect optimization time from at least $\sim$8–9 h (HSE06) to $\sim$1–2 min while retaining near-DFT accuracy, transforming comprehensive, composition- and charge-resolved defect surveys from intractable to routine. We have deployed this framework as part of a Jupyter notebook-based nanoHUB tool which will allow users to upload CIF files of Cd/Zn-Te/Se/S structures, auto-generate relevant defects or complexes, and compute their defect formation energies as functions of Fermi level and chemical potentials conditions, bypassing expensive first principles workflows. Together, these advances provide a scalable, charge-aware pathway to map defect landscapes in chemistries relevant to CdSeTe solar cell devices and beyond, accelerating the dopant/process optimization and ultimately closing the voltage deficit in this important thin-film photovoltaic platform.